# Misspecification uncertainties in near-deterministic regression

### v2

@@ Abstract
The expected loss is an upper bound to the model generalization error which admits robust PAC-Bayes bounds for learning. However, loss minimization is known to ignore misspecification, where models cannot exactly reproduce observations. This leads to significant underestimates of parameter uncertainties in the large data, or underparameterized, limit. We analyze the generalization error of near-deterministic, misspecified and underparametrized surrogate models, a regime of broad relevance in science and engineering. We show posterior distributions must cover every training point to avoid a divergent generalization error and derive an ensemble ansatz that respects this constraint, which for linear models incurs minimal overhead. The efficient approach is demonstrated on model problems before application to high dimensional datasets in atomistic machine learning. Parameter uncertainties from misspecification survive in the underparametrized limit, giving accurate prediction and bounding of test errors.
I Introduction
@@
Surrogate models are widely used across science and engineering, notably to efficiently approximate the action of computationally intensive simulation engines  (Alizadeh et al., 2020; Deringer et al., 2019; Lapointe et al., 2020; Nyshadham et al., 2019; Montes de Oca Zapiain et al., 2022; Bonatti et al., 2022; Kudela and Matousek, 2022). Domain expertise is leveraged to design optimal features and model architecture before parameter inference is carried out based on training data generated by the engines. In this paper, we consider the common scenario where the learning problems share three key characteristics:
@@
  • Simulation engines are near deterministic: simulation outputs have vanishing aleatoric uncertainty, e.g. the atomic energy $V({\bf X})$ is a deterministic function of atomic positions ${\bf X}$.
@@
  • Surrogate models are misspecified: no single choice of the $P$ model parameters can match all $N$ training observations, such that model averaging is required to capture uncertainties.
@@
  • Surrogate models are underparametrized: large quantities of training data are available, such that the standard posterior from Bayesian inference gives vanishing parameter uncertainties. This typically occurs in the limit $P/N\ll 1$ for finite capacity models.
Misspecification affects both finite capacity models and deep learning approaches with finite training resources  (Lahlou et al., 2021; Psaros et al., 2023). A misspecification-aware learning scheme should minimize the cross entropy between predicted and observed data distributions, known as the generalization error, or population risk (see (Alquier, 2021; Morningstar et al., 2022)). However, this is typically not tractable for learning, both due to numerical instabilities and the lack of theoretical bounds for convergence.
@@
Bayesian learning schemes instead minimize the expected loss (log likelihood), an upper bound to the generalization error, which is known to ignore misspecification but admits a robust learning scheme. Epistemic uncertainties are thus severely underestimated, vanishing in the underparametrized limit, with broad implications for surrogate model selection, uncertainty quantification and error propagation, as noted by multiple groups(Masegosa, 2020; Kato et al., 2022; Lotfi et al., 2022; Lahlou et al., 2021; Psaros et al., 2023; Imbalzano et al., 2021).
@@
We analyze the generalization error of deterministic, misspecified and under-parametrized surrogate models, i.e. those typically used to approximate simulation engines in science and engineering. Our focus is the underparametrized regime, meaning we do not focus on epistemic errors important for finite $N$. As posterior distributions become unbounded in the deterministic limit, this is only taken as a theoretical device to derive conditions that any posterior must obey to avoid a divergent generalization error. We then design an ensemble ansatz that respects this condition and admits a robust learning scheme at finite $N$, which reduces to a constrained loss minimization problem for large $N/P$. Our main contributions are:
@@
  • We define a pointwise optimal parameter sets (POPS) for each training point, for which model predictions are exact. Parameter distributions must have mass in every POPS to avoid a divergent generalization error. An ensemble ansatz is proposed that respects POPS occupancy and gives a finite generalization error.
@@
  • For linear models, the variational minimum of our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer. This is our main result, an efficient scheme to incorporate misspecification into linear regression with minimal computational overhead.
  • We compare our variational minimum to numerial optimization of our ensemble ansatz and show how our scheme provides robust bounds on test errors. This is confirmed in numerical experiments on challenging datasets from atomic machine learning.
@@
(Germain et al., 2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis for regression problems with unbounded losses, which bounds true expectations by empirical likelihood expectations through concentration inequalities. This provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization. In particular, minimizing the PAC-Bayes bound for the expected loss gives the familiar relation between posterior, prior and likelihood from Bayesian inference.
@@
In recent years, multiple groups have used the PAC-Bayes framework to derive tighter bounds for the generalization error than that provided by the expected loss  (Masegosa, 2020; Germain et al., 2016; Lahlou et al., 2021; Psaros et al., 2023; Lotfi et al., 2022; Morningstar et al., 2022). However, these efforts focus on probabilistic regression settings, often with neural network models far from the underparametrized limit, meaning aleatoric, epistemic and misspecification errors must be considered jointly.
@@
(Masegosa, 2020) derived second order PAC-Bayes bounds for the generalization error, minimized through a variational or ensemble ansatz. (Morningstar et al., 2022) developed estimators to assess disagreement between the generalization error and expected loss, deriving specialized PAC bounds for theoretical guarantees. (Lahlou et al., 2021) considered misspecification of neural networks, jointly training a minimum loss surrogate and an independent predictor of misspecification error. (Lotfi et al., 2022) considered misspecification in the context of deep model selection, deriving conditional likelihoods better aligned with the generalization error. As various definitions for the distinction between epistemic and misspecification uncertainties exist (Kato et al., 2022), we define misspecification as non-aleatoric uncertainties which survive in the underparametrized limit, where the minimizer of PAC-Bayes bounds for the expected loss, which coincides with the posterior from Bayesian inference, predict vanishing parameter uncertainties. This definition is made more precise in section II.3. Our epistemic uncertainties thus coincide with standard estimates from loss minimization.
We summarize known results for the generalization error and expected loss, PAC-Bayes bounds for the regression loss, and the sub-optimality of loss minimization for misspecified problems in section II. Section III contains our central theoretical contribution, an analysis of the generalization error for near-deterministic surrogate models and a simple ensemble ansatz which can be efficiently evaluated for linear models. Section IV applies our ansatz to high-dimensional linear regression problems, using synthetic data and challenging applications in atomic machine learning.
II Bayesian surrogate models
II.1 Deterministic Simulation Engines
@@
II.2 Deterministic Surrogate Models and the Generalization Error
@@
Unlike typical Bayesian schemes, we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of the quasi-deterministic simulation engine ${\mathcal{E}}$. This has implications for model selection criteria, as we discuss in appendix B.
@@
II.3 The expected loss and PAC estimation
PAC-Bayes analysis (Alquier, 2021) provides concentration inequalities (Hoeffding, 1994) that bound expectations (3) over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$ with some probability $1-\xi$. Whilst originally designed for classification problems where the loss is bounded to $[0,1]$, the treatment can be extended to regression settings with unbounded losses, as discussed in (Germain et al., 2016) and appendix A. In this setting, available PAC-Bayes bounds are for $\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{\mathcal{D}}$, where $f({\mathbf{Y}},{\mathbf{X}})$ to be bounded (i.e. $\epsilon>0$) and can expressed as an expectation over $\pi({\bm{\Theta}})$, i.e. $\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{\mathcal{D}}=\int_{\mathcal{P}}\langle f({\mathbf{Y}},{\mathbf{X}},{\bm{\Theta}})\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}$. As ${\mathcal{G}}[\pi]$ is not in this form, the bound used in PAC-Bayesian schemes is for the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood. The Jensen inequality for convex functions $f(\langle x\rangle)\leq\langle f(x)\rangle$ gives, with $f(x)=-\ln|x|$,
$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$
As discussed by (Germain et al., 2016) the expected regression loss ${\mathcal{L}}[\pi]$ admits the PAC-Bayes upper bound, holding with probability $1-\xi$, of
${\mathcal{L}}[\pi]$$\leq\hat{{\mathcal{L}}}_{N}[\pi]+\frac{1}{N}\left(KL(\pi|\pi_{0})-\ln|\xi|\right)+C_{0}[\pi_{0}],$
^$\hat{{\mathcal{L}}}_{N}[\pi]$ $\equiv\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}}){\rm d}\pi({\bm{\Theta}}),$ (7)
where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, $KL(\pi|\pi_{0})=\int_{\mathcal{P}}\ln|\pi/\pi_{0}|{\rm d}\pi({\bm{\Theta}})$ is the cross entropy between $\pi$ and some prior $\pi_{0}({\bm{\Theta}})$ and $C_{0}[\pi_{0}]$ is constant with respect to $\pi$ and has a finite upper bound discussed in appendix A. The minimizer of the PAC-Bayes upper bound (7) is that of ^$\hat{{\mathcal{L}}}_{N}[\pi]+KL(\pi|\pi_{0})/N$, giving the posterior from Bayesian inference (Germain et al., 2016)-
@@ II.3 The expected loss and PAC estimation
Appendix B discusses this bound in the context of Bayesian model selection. Using Laplace’s method (appendix C) as $N/P\to\infty$ gives $\hat{L}_{N}({\bm{\Theta}})\to L({\bm{\Theta}})$, with
@@
[…]
$L({\bm{\Theta}})=-\langle\ln|\mathcal{N}({\mathcal{E}}({\mathbf{X}})|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}),{\bm{\Sigma}}_{\mathcal{Y}})|\rangle_{\mathcal{X}},$ (10)
[…]
@@
${\bm{\Sigma}}^{*}_{\mathcal{L}}=\langle[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})][{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]^{\top}\rangle$ is the error covariance around ${\bm{\Theta}}^{*}_{\mathcal{L}}$ and we use ${\rm Tr}({\bm{\Sigma}}_{\mathcal{Y}})=\mathcal{O}(\epsilon^{2})$. As noted by (Masegosa, 2020), it is clear that the maximum likelihood parameter distribution $\pi_{\mathcal{L}}({\bm{\Theta}})$ is not a minimizer of the generalization error under misspecification.
We emphasize that the primary motivation of this section is to show that whilst the generalization error is the central quantity of interest for learning under misspecification, we typically only have access to an upper bound in the form of a PAC-Bayes concentration inequality (7) for the expected loss. Under model misspecification it is clear that minimization of this upper bound in sub-optimal, in that the generalization error diverges as $1/\epsilon^{2}$. In the next section we analyze the generalization error as $\epsilon\to 0$ to derive a condition for $\pi({\bm{\Theta}})$ that ensures a finite generalization error. We then design an ensemble ansatz which respects this condition and can be parametrized through a variational minimization of the PAC-Bayes upper bound (7).
III Analysis of the generalization error in the near-deterministic limit
We consider the generalization error (6) in the near-deterministic limit, ${\mathcal{G}}_{0}[\pi]\equiv\lim_{\epsilon\to 0}{\mathcal{G}}[\pi]$. The integral over ${\mathbf{Y}}$ in (6) concentrates at ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$; using $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$ yields, using Laplace’s method

### v3

# Parameter uncertainties for imperfect surrogate models in the low-noise regime

@@ Abstract
Bayesian regression determines model parameters by minimizing the expected loss, an upper bound to the true generalization error. However, the loss ignores misspecification, where models are imperfect. Parameter uncertainties from Bayesian regression are thus significantly underestimated and vanish in the large data limit. This is particularly problematic when building models of low-noise, or near-deterministic, calculations, as the main source of uncertainty is neglected. We analyze the generalization error of misspecified, near-deterministic surrogate models, a regime of broad relevance in science and engineering. We show posterior distributions must cover every training point to avoid a divergent generalization error and design an ansatz that respects this constraint, which for linear models incurs minimal overhead. This is demonstrated on model problems before application to thousand dimensional datasets in atomistic machine learning. Our efficient misspecification-aware scheme gives accurate prediction and bounding of test errors where existing schemes fail, allowing this important source of uncertainty to be incorporated in computational workflows.
@@
Surrogate models are widely used across science and engineering to efficiently approximate the action of computationally intensive simulation engines  Alizadeh et al. (2020); Deringer et al. (2019); Lapointe et al. (2020); Nyshadham et al. (2019); Montes de Oca Zapiain et al. (2022); Bonatti et al. (2022); Kudela and Matousek (2022). Domain expertise is leveraged to design optimal features and model architecture before parameters are inferred from training data generated by the engines. In a broad range of applications, the inference problem shares three key characteristics:
@@
a) Simulation engines are near deterministic: outputs have vanishing aleatoric uncertainty, e.g. the atomic energy $V({\bf X})$ is deterministic to atomic positions ${\bf X}$.
@@
b) Surrogate models are misspecified: no single choice of the $P$ parameters can match all $N$ observations, meaning parameters are intrinsically uncertain.
@@
c) Surrogate models are underparametrized: large quantities of training data are available, such that standard epistemic uncertainty estimates from Bayesian inference vanish, typically corresponding to the regime $P/N\ll 1$.
@@
Misspecification affects both finite capacity models and deep learning approaches with finite training resources  Lahlou et al. (2021); Psaros et al. (2023). A misspecification-aware learning scheme should minimize the cross entropy between predicted and observed data distributions, known as the generalization error, or under certain technical definitions, the population risk (see Alquier (2021); Morningstar et al. (2022)). However, the generalization error is typically not tractable for learning, both due to numerical instabilities and the lack of theoretical bounds for convergence. Bayesian learning schemes instead minimize an upper bound, the expected loss (log likelihood), which ignores misspecification but admits a robust learning scheme. Epistemic uncertainties are thus severely underestimated, vanishing in the underparametrized limit, with broad implications for surrogate model selection, uncertainty quantification and error propagation, as noted by multiple groups Masegosa (2020); Kato et al. (2022); Lotfi et al. (2022); Lahlou et al. (2021); Psaros et al. (2023); Imbalzano et al. (2021).
@@
In this paper, we analyze the generalization error of deterministic, misspecified and under-parametrized surrogate models, i.e. those typically used to approximate simulation engines in science and engineering. As posterior distributions strictly become unbounded in the deterministic limit, this is only taken as a theoretical device to derive conditions that any posterior must obey to avoid a divergent generalization error. We then design an ansatz that respect this condition and admits a robust learning scheme at finite $N$, with standard treatment of epistemic ($N$-dependent) errors, reducing to a constrained loss minimization problem for large $N/P$. Our final misspecification-aware parameter posterior has finite uncertainty even as $N/P\to\infty$ with a provably lower generalization error than the Gibbs posterior of Bayesian inference.
Our main contributions are:
@@
a) We define pointwise optimal parameter sets (POPS) for each training point, within which model predictions are exact. Parameter distributions must have mass in every POPS to avoid a divergent generalization error. b) We use the ensemble of loss minimizers from each POPS to design two ansatz posteriors, a reweighted $\mathcal{O}(N)$ ensemble and a bounding $\mathcal{O}(P)$ hypercube, both of which respect POPS occupancy and give a finite generalization error, lower than the Gibbs posterior from inference.
@@
c) For linear models, our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer. This gives an efficient scheme to incorporate misspecification into linear regression with minimal computational overhead.
d) We test the approach on synthetic datasets, showing how our scheme provides robust bounds on test errors. This performance is maintained under application to thousand-dimensional datasets from atomic machine learning, untreatable with existing methods.
.1 Related work
@@
Germain and coworkers Germain et al. (2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis for regression problems with unbounded losses, which bounds true expectations by empirical likelihood expectations through concentration inequalities. This provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization. In particular, minimizing the PAC-Bayes bound for the expected loss gives the familiar relation between posterior, prior and likelihood from Bayesian inference.
@@
In recent years, multiple groups have used the PAC-Bayes framework to derive tighter bounds for the generalization error than that provided by the expected loss  Masegosa (2020); Germain et al. (2016); Lahlou et al. (2021); Psaros et al. (2023); Lotfi et al. (2022); Morningstar et al. (2022). However, these efforts focus on probabilistic regression settings, often with neural network models far from the underparametrized limit, meaning aleatoric, epistemic and misspecification errors must be considered jointly.
@@
Masegosa Masegosa (2020) derived second order PAC-Bayes bounds for the generalization error minimized through a variational or ensemble ansatz. Morningstar et al. Morningstar et al. (2022) developed estimators to assess disagreement between the generalization error and expected loss, deriving specialized PAC bounds for theoretical guarantees. Lahlou et al. Lahlou et al. (2021) considered misspecification of neural networks, jointly training a minimum loss surrogate and an independent predictor of misspecification error. Lofti et al. Lotfi et al. (2022) considered misspecification in the context of deep model selection, deriving conditional likelihoods better aligned with the generalization error.
As various definitions for the distinction between epistemic and misspecification uncertainties exist Kato et al. (2022), we define misspecification as non-aleatoric uncertainties which survive in the underparametrized limit, where the minimizer of PAC-Bayes bounds for the expected loss, which coincides with the posterior from Bayesian inference, predict vanishing parameter uncertainties. This definition is made more precise in section I.4. Our epistemic uncertainties thus coincide with standard estimates from loss minimization.
I Bayesian surrogate models
I.1 Deterministic simulation engines
@@
I.2 Deterministic surrogate models
@@
Unlike typical Bayesian schemes, we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of ${\mathcal{E}}$. This has implications for model selection criteria, discussed in the supplementary material (SM) Swinburne and Perez (2024).
I.3 The generalization error and expected loss
@@
In principle any learning scheme should aim to minimize the non-convex functional ${\mathcal{G}}[\pi]$, but in practice learning schemes target a convex upper bound, the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood. Using the Jensen inequality $-\langle\ln x\rangle\leq-\ln\langle x\rangle$ gives
$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (7)
${\mathcal{L}}[\pi]$ is clearly minimized by a delta function $\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$, where ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{\mathcal{P}}\left\langle-\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}$, i.e. vanishing parameter uncertainty. The next section shows how this emerges from the inference posterior at finite $N/P$.
I.4 PAC-Bayes estimation and connection to inference
PAC-Bayes analysis Alquier (2021) provides concentration inequalities Hoeffding (1994) that bound expectations over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$ with some probability. Whilst originally designed for classification problems, the approach has been extended to regression settings with unbounded losses Germain et al. (2016). As discussed in the SM Swinburne and Perez (2024), The expected loss ${\mathcal{L}}[\pi]$ admits the PAC-Bayes upper bound, holding with probability $1-\xi$, of
${\mathcal{L}}[\pi]\leq\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi}{\pi_{0}}\right|{\rm d}\pi({\bm{\Theta}})-\frac{\ln|\xi|}{N}+C_{0},$ (8)
where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, $\pi_{0}({\bm{\Theta}})$ is some prior and $C_{0}$ is a constant Swinburne and Perez (2024). The minimizer of the upper bound (21) is the posterior from Bayesian inference Germain et al. (2016)
@@ I.4 PAC-Bayes estimation and connection to inference
As $N/P\to\infty$ the posterior is strongly peaked around a loss minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{{\bm{\Theta}}\in{\mathcal{P}}}L({\bm{\Theta}})$. Defining ${\bm{\Sigma}}^{-1}_{0}$ as the curvature of $-\ln\pi_{0}({\bm{\Theta}})$ around ${\bm{\Theta}}^{*}_{\mathcal{L}}$, we use Laplace’s method Swinburne and Perez (2024) as $N/P\to\infty$ to write
@@
${\bm{\Sigma}}^{*}_{\mathcal{L}}=\langle[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})][{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]^{\top}\rangle$ is the error covariance around ${\bm{\Theta}}^{*}_{\mathcal{L}}$ and we use ${\rm Tr}({\bm{\Sigma}}_{\mathcal{Y}})=\mathcal{O}(\epsilon^{2})$. Under misspecification it is thus clear that minimizing this upper bound is sub-optimal Masegosa (2020) as the generalization error diverges as $1/\epsilon^{2}$. In the next section we analyze the generalization error as $\epsilon\to 0$ to derive a condition for any $\pi({\bm{\Theta}})$ that ensures a finite generalization error.
II Analysis of the generalization error in the near-deterministic limit
We consider the generalization error (6) in the near-deterministic limit, ${\mathcal{G}}_{0}[\pi]\equiv\lim_{\epsilon\to 0}{\mathcal{G}}[\pi]$. The integral over ${\mathbf{Y}}$ in (6) concentrates at ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$; using $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$ yields, using Laplace’s method Swinburne and Perez (2024)